# VINP: Variational Bayesian Inference with Neural Speech Prior for Joint ASR-Effective Speech Dereverberation and Blind RIR Identification

### v2

@@ III-B1 E-step
whose precision and mean have close-form solutions as
@@
where is a smoothing factor, and are the estimates from the previous VEM iteration.
The prior of anechoic speech narrows down the infinite number of possible solutions when decoupling the anechoic speech and the RIR, which is the key to achieving dereverberation [42]. The mean of the posterior distribution is the MAP estimate of anechoic spectrum . It is also worth noticing that the E-step can be implemented in parallel across all T-F bins.
@@ III-B2 M-step
In this step, VEM updates the noise precision and CTF filter by maximizing the logarithmic joint probability of the anechoic speech and observation, which is
@@ III-B3 Initialization of VEM Parameters
The initialization of parameters plays a crucial role in the convergence of VEM. Before iteration, the mean and variance of the anechoic speech posterior are set to zero and the power spectrum of reverberant recording, respectively. This means we have
The CTF coefficients in each frequency band are set to 0 except that the first coefficient is set to 1, which means
Because even during speech activity, the short-term power spectral density of observation often decays to values that are representative of the noise power level [43], the initial variance of the additional noise is set to the minimum power in each frequency band, which means
The VBI procedure is summarized in Algorithm 1. The final outputs of VBI procedure are the complex-valued anechoic spectrum and CTF filter estimates. Notice that in the proposed framework, the prior distribution of anechoic speech is not updated during iterations. Therefore, the DNN only infers once.
@@ III-C Transformation to Waveforms
After the VBI procedure, both the anechoic spectrum and the CTF filter are estimated. We need to further transform these T-F representations into waveforms. By applying inverse STFT to the anechoic spectrum estimate, we can easily get the anechoic speech waveform. For the estimation of RIR, we design a pseudo intrusive measurement process as follows.
A common method for intrusive measuring the RIR of an acoustical system is to apply a known excitation signal and measure the microphone recording [44]. Playing an excitation signal with a loudspeaker, the noiseless microphone recording can be written as
@@
where is the initial radian frequency and is the final radian frequency of the sweep with duration . Through an ideal inverse filter , the excitation signal can be transformed into a Dirac’s delta function , as
For the logarithmic sine sweep excitation, the inverse filter is an amplitude-modulated and time-reversed version of itself [44, 45]. The RIR can be estimated by convolving the measurement with the inverse filter as
In our approach, the excitation signal is convoluted by the CTF estimates (along the time dimension) to get a pseudo measurement in the T-F domain as
where , and are the STFT coefficient of and , respectively. Finally, after applying inverse STFT to , we use the pseudo measurement to estimate the RIR waveform by
@@ IV-A Datasets
VINP is designed for both speech dereverberation and blind RIR identification. These two tasks share the same DNN for speech prior distribution prediction. Therefore, we use a single training set and two different test sets.
@@ IV-A1 Training Set
The training set is composed of 200 hours of high-quality English speech utterances from the corpora of EARS [46], DNS Challenge [47], and VCTK [48]. All speech utterances are downsampled to 16 kHz if necessary. We simulate 100,000 pairs of reverberant and direct-path RIRs using the gpuRIR toolbox [49]. The simulated speaker and microphone are randomly placed in rooms with dimensions randomly selected within a range of 3 m to 15 m (for length and width) and 2.5 m to 6 m (for height). The minimum distance between the speaker/microphone and the wall is 1 m. Reverberant RIRs have RT60 values uniformly selected within the range of 0.2 s to 1.5 s. Direct-path RIRs are generated using the same geometric parameters as the reverberant ones but with an absorption coefficient of 0.99. Noise recordings from the Noise92 corpus and the training set from the REVERB Challenge [50] are employed. The signal-to-noise ratio (SNR) is randomly selected within the range of 5 dB to 20 dB.
@@ IV-A2 Test Set for Speech Dereverberation
For speech dereverberation, we utilize the official single-channel test set from the REVERB Challenge [50], which includes both simulated data (marked as ’SimData’) and real recordings (marked as ’RealData’).
In SimData, there exist six distinct reverberation conditions: three room volumes (small, medium, and large), and two distances between the speaker and the microphone (50 cm and 200 cm). The RT60 values are approximately 0.25 s, 0.5 s, and 0.7 s. The noise is stationary background noise, mainly generated by air conditioning systems. SimData has a SNR of 20 dB.
@@ IV-A3 Test Set for Blind RIR Identification
A test set named ’SimACE’ is constructed to evaluate blind RIR estimation. In SimACE, microphone signals are simulated by convolving the clean speech from the ’si_et_05’ subset in WSJ0 corpus [51] with the downsampled recorded RIRs from the ’Single’ subset in ACE Challenge [52], and adding noise from the test set in REVERB Challenge [50]. The minimum and maximum RT60s are 0.332 s and 1.22 s, respectively. More details about the RIRs can be found in [52]. We create the SimACE test set because the RT60 labels of the RIRs in SimACE are more accurate than those in REVERB Challenge. SimACE has a SNR of 20 dB.
@@ IV-B1 Data Representation
Before feeding the speech into VINP, the reverberant waveform is normalized by its maximum absolute value. Subsequently, the utterance is transformed using STFT with a Hann window of 512 samples (32 ms) and a hop length of 128 samples (8 ms).
@@ IV-B2 DNN Architecture
VINP is able to employ any discriminative dereverberation DNNs to predict the prior distribution of anechoic speech. In this paper, we evaluate with the network proposed in TCN+SA+S [15] and a modified network of the Mamba version of oSpatialNet [53], resulting in two different versions marked as ’VINP-TCN+SA+S’ and ’VINP-oSpatialNet’, respectively. For both versions, the input and output of DNNs are the 10-based logarithmic magnitude spectra. To avoid numerical issues, a small constant is added to the magnitude spectra before taking the logarithm.
The DNN architecture in VINP-TCN+SA+S is composed of temporal convolutional networks (TCNs) and self-attention layers, as shown in Fig. 3. We use the same settings as the original TCN+SA+S [15], except that there is no activation function after the output layer and we do not use dropout.
The DNN architecture in VINP-oSpatialNet is derived from oSpatialNet [53], as depicted in Fig. 3. A two-dimensional convolution layer with a kernel size of is employed to expand the input to 96 dimensions. The remaining modules are the same as those in the original oSpatialNet. Except that, for offline processing, we replace the second forward Mamba layer in the narrow-band block with a backward Mamba layer by simply reversing the input and output along the time dimension.
@@ IV-B3 Training Configuration
For DNN training, we set the hyperparameter in the loss function to . The speech utterances are segmented into 3 s. The batch size of VINP+TCN+SA+S and VINP-oSpatialNet are set to 16 and 4, respectively. The AdamW optimizer [54] with an initial learning rate of 0.001 is employed. The learning rate exponentially decays with and in VINP-TCN+SA+S and VINP-oSpatialNet, respectively. Gradient clipping is applied with a L2-norm threshold of 10. The training is carried out for 800,000 steps in total. We average the model weights of the best three epochs as the final model.
@@ IV-B4 VBI Settings
The length of the CTF filter is set to . The fixed smoothing factor in Eq. (22) is set to 0.7 to obtain a stable result. In rare cases, if the likelihood of the complete data decreases, the VBI iteration will be stopped prematurely. Since the fundamental frequency of human speech is always higher than 85 Hz [55], we ignore the three lowest frequency bands and set their coefficients to zero to avoid the effect of extremely low SNR in these bands. Therefore, VBI processes a total of 254 frequency bands.
@@ IV-B6 RT60 and DRR Estimation
RT60 and DRR are key acoustic parameters that feature the characteristic of RIR, which are thus used in this work for evaluating the accuracy of RIR identification. RT60 and DRR can be directly calculated from RIR.
@@
Since the fact that sound energy decays exponentially over time, a linear fitting is applied to a portion of logarithmic EDC, and the slope is utilized to calculate RT60. During this process, the key to linear fitting lies in the heuristic strategy of selecting the fitting range. In this work, the starting point of linear fitting is selected within the range of a time delay from 20 ms to 50 ms at the direct-path sample, which represents the sample corresponding to the maximum value of the RIR. Meanwhile, the ending point is established at a 5 dB attenuation relative to the starting point. We fit all intervals that meet the conditions and apply the result with the maximum absolute Pearson correlation coefficient to the estimation process of RT60 as
@@
DRR refers to the ratio between the energy of the direct-path sound and the reverberant sound. In this work, the DRR is defined as
where the direct-path signal arrives at the th sample, and is the additional sample spread for the direct-path response which typically corresponds to 2.5 ms [52].
@@ IV-C1 Speech Dereverberation
We compare VINP with various advanced dereverberation approaches including GWPE [4], SkipConvNet [57]222https://github.com/zehuachenImperial/SkipConvNet, CMGAN [23]333https://github.com/ruizhecao96/CMGAN, and StoRM [25]444https://github.com/sp-uhh/storm. In addition, to demonstrate the effect of VINP, the DNNs in VINP-TCN+SA+S and VINP-oSpatialNet are also compared, i.e. TCN+SA+S [15] and oSpatialNet [53]. In TCN+SA+S, we use the recommended Griffin-Lim’s iterative algorithm [58] to restore the phase spectrum. In oSpatialNet, for a fair comparison, the backward Mamba is also applied to oSpatialNet. Different from VINP-oSpatialNet, the oSpatialNet employed for comparison processes the complex-valued speech spectrum as the original paper [53]. And this version is marked as ’oSpatialNet*’.
For all comparison methods, we use their recommended configuration and official codes (if available). All DL-based approaches are trained from scratch on the same training set. GWPE is implemented using the NaraWPE python package [59]555https://github.com/fgnt/nara_wpe.
@@
VINP-TCN+SA+S | 4.7 | 0.7+0.27iterations |
VINP-oSpatialNet | 1.7 | 36.6+0.27iterations |
@@ IV-C2 Blind RIR Identification
We measure RT60 and DRR through the RIR estimates to show the effectiveness of RIR identification. The comparison methods include several classical blind RT60 and DRR estimation methods, such as Ratnam’s approach [60]666https://github.com/nuniz/blind_rt60 and Jeub’s approach [61]777https://ww2.mathworks.cn/matlabcentral/fileexchange/32752-blind-direct-to-reverberant-energy-ratio-drr-estimation. As a DL-based blind RIR identification method, BUDDy [35]888https://github.com/sp-uhh/buddy is also compared. All these methods are implemented through the open-sourced official codes. Specifically, because the output of BUDDy is the RIR waveform, we employ the same implementation as in VINP to estimate the RT60 and DRR. Notice that the acoustic conditions of our training set and test set are mismatched. BUDDy, as an unsupervised approach, is designed to bridge the performance gap between scenarios with matched and mismatched acoustic conditions. Therefore, we directly employ the model parameters provided by the authors without retraining.
Method | SimData | RealData | ||||||||||
13 | ||||||||||||
PESQ | ESTOI | DNSMOS | WER (%) | DNSMOS | WER (%) | |||||||
13 | ||||||||||||
P.835 | P.808 | tiny | small | medium | P.835 | P.808 | tiny | small | medium | |||
Unprocessed | 1.48 | 0.70 | 2.37 | 3.20 | 13.1 | 5.6 | 4.6 | 1.31 | 2.82 | 24.1 | 7.9 | 5.7 |
Oracle | - | - | 3.76 | 3.90 | 7.6 | 4.5 | 4.0 | - | - | - | - | - |
GWPE [4] | 1.55 | 0.72 | 2.41 | 3.22 | 12.1 | 5.5 | 4.6 | 1.42 | 2.83 | 21.4 | 6.8 | 5.7 |
SkipConvNet [57] | 2.12 | 0.81 | 3.20 | 3.60 | 13.3 | 6.3 | 5.2 | 2.84 | 3.32 | 24.5 | 9.3 | 7.3 |
CMGAN [23] | 2.85 | 0.90 | 3.82 | 3.81 | 9.5 | 5.1 | 4.4 | 3.87 | 4.00 | 12.9 | 5.9 | 5.0 |
StoRM [25] | 2.34 | 0.86 | 3.73 | 3.96 | 11.4 | 6.2 | 5.1 | 3.72 | 4.01 | 17.5 | 10.2 | 8.0 |
[…]
oSpatialNet* [53] | 2.87 | 0.92 | 3.57 | 3.88 | 8.9 | 4.9 | 4.3 | 3.48 | 3.87 | 10.5 | 5.4 | 4.5 |
VINP-TCN+SA+S (prop.) | 2.52 | 0.87 | 3.47 | 3.88 | 8.9 | 5.1 | 4.3 | 3.18 | 3.77 | 11.6 | 6.1 | 5.3 |
VINP-oSpatialNet (prop.) | 2.82 | 0.90 | 3.49 | 3.86 | 8.3 | 4.8 | 4.1 | 3.33 | 3.80 | 8.9 | 5.0 | 4.3 |

### v3

@@ III-B1 E-step
whose precision and mean have closed-form solutions as
@@
where is a smoothing factor, and are the estimates from the previous VEM iteration. The mean of the posterior distribution is the MAP estimate of anechoic spectrum . It is also worth noticing that the E-step can be implemented in parallel across all T-F bins.
@@ III-B2 M-step
In this step, VEM updates the noise precision and CTF filter by maximizing the expected logarithmic likelihood of the complete data, which is
@@ III-B3 Initialization of VEM Parameters
The initialization of parameters plays a crucial role in the convergence of VEM. We use an uninformative initialization in VINP. Before iteration, the mean and variance of the anechoic speech posterior are set to zero and the spectrogram of reverberant recording respectively as
The CTF coefficients in each frequency band are set to 0, except that the first coefficient is set to 1, which means
Because even during speech activity, the short-term power spectral density of observation often decays to values representative of the noise power level [47], the initial variance of the additional noise in each frequency band is set to the minimum power across all frames, which means
III-B4 Outputs of VBI
During VEM iteration, in each frequency band, there is a narrow-band expected logarithmic likelihood of complete data, written as . Once VEM reaches the preset maximum number of iterations, it terminates. In each frequency band, the output of VBI procedure is the estimated complex-valued clean spectrum and CTF filter corresponding to the maximum expected logarithmic likelihood.
The VBI procedure is summarized in Algorithm 1. The prior distribution of anechoic speech serves to constrain the infinite set of potential solutions inherent in the signal model, which is the key to effective dereverberation [48]. Note that in the proposed framework, the prior distribution of anechoic speech remains unchanged during iterations. Therefore, the DNN performs inference through a single forward pass without iterative optimization.
@@ III-C Transformation to Waveforms
After the VBI procedure, both the anechoic spectrum and the CTF filter are estimated. We need to further transform these T-F representations into waveforms. By applying inverse STFT to the anechoic spectrum, we can easily obtain the anechoic speech waveform. However, CTF is the model of reverberation in the T-F domain rather than the STFT of RIR. Therefore, we cannot convert the CTF filter to RIR directly. In VINP, we design a pseudo intrusive measurement process to obtain the RIR estimate as follows.
Reviewing existing approaches, a common method for intrusive measuring the RIR of an acoustical system is to apply a known excitation signal and measure the microphone recording [49]. When a loudspeaker plays an excitation signal , the noiseless microphone recording can be written as
@@
where is the initial radian frequency and is the final radian frequency of the sweep with duration . Through an ideal inverse filter , the excitation signal can be transformed into an impulse , as
For the logarithmic sine sweep excitation, the inverse filter is an amplitude-modulated and time-reversed version of itself [49, 50]. Subsequently, the RIR can be estimated by convolving the measurement with the inverse filter as
In our approach, the excitation signal is convoluted with the CTF filter in the T-F domain along the time dimension to yield a pseudo measurement spectrum as
where , and are the STFT coefficient of and , respectively. Applying inverse STFT to , we use the pseudo measurement and the inverse filter to estimate the RIR waveform as
@@ IV-A Datasets
VINP is designed for joint speech dereverberation and blind RIR identification. We use a single training set and two different test sets. To ensure sampling rate consistency, all audio utterances in these datasets are downsampled to 16 kHz.
@@ IV-A1 Training Set
The training utterances are generated by convolving anechoic source speech with reverberant and direct-path RIRs, followed by noise addition. The source speech consists of 200 hours of high-quality English speech utterances, including the clean speech from DNS Challenge [51] and VCTK [52] that have a raw DNSMOS p.835 score [53] exceeding 3.5, as well as the whole EARS [54] corpus. We simulate 100,000 pairs of reverberant and direct-path RIRs using the gpuRIR toolbox [55]. The simulated speaker and omnidirectional microphone are randomly placed in rooms with dimensions randomly selected within a range of 3 m to 15 m for length and width, and 2.5 m to 6 m for height. The minimum distance between the speaker/microphone and the wall is 1 m. Reverberant RIRs have RT60s uniformly distributed within the range of 0.2 s to 1.5 s. Direct-path RIRs are generated using the same geometric parameters as the reverberant ones but with an absorption coefficient of 0.99. Noise recordings from NOISEX-92 [56] and the training set of REVERB Challenge [57] are used. The signal-to-noise ratio (SNR) is uniformly distributed within the range of 5 dB to 20 dB.
@@ IV-A2 Test Set for Speech Dereverberation
For speech dereverberation, we utilize the official single-channel test set from the REVERB Challenge [57], which includes both simulated recordings (marked as ’SimData’) and real recordings (marked as ’RealData’).
In SimData, there exist six distinct reverberation conditions: three room volumes (small, medium, and large), and two distances between the speaker and the microphone (50 cm and 200 cm). The RT60 values are approximately 0.25 s, 0.5 s, and 0.7 s. The noise is stationary, mainly generated by air conditioning systems. SimData has a SNR of 20 dB.
@@ IV-A3 Test Set for Blind RIR Identification
A test set named ’SimACE’ is constructed to evaluate blind RIR identification. In SimACE, microphone signals are simulated by convolving the clean speech from the ’si_et_05’ subset in WSJ0 corpus [58] with the downsampled recorded RIRs from the ’Single’ subset in ACE Challenge [59], and adding noise from the test set in REVERB Challenge [57]. The minimum and maximum RT60s are 0.332 s and 1.22 s, respectively. More details about the RIRs can be found in [59]. SimACE has a SNR of 20 dB.
Notice that, for both tasks, the RIRs in the test sets are real-measured instead of simulated, which may lead to some mismatch with the training set.
@@ IV-B1 Data Representation
Before feeding the speech into VINP, the reverberant waveform is normalized by its maximum absolute value. The STFT analysis window and synthesis window are set to Hann windows with a length of 512 samples and 75% overlap.
@@ IV-B2 DNN Architecture
VINP is able to employ any discriminative dereverberation DNNs to estimate the prior distribution of anechoic speech. Two representative dereverberation backbones TCN+SA+S [18] and oSpatialNet-Mamba [60] are selected. Among those backbones, TCN+SA+S includes a self-attention module to produce dynamic representations given input features, a temporal convolutional network (TCN) to learn a nonlinear mapping from such representations to the magnitude spectrum of anechoic speech, and a 1-D convolution module to smooth the enhanced magnitude among adjacent frames. Different from TCN+SA+S, oSpatialNet-Mamba employs a 1-D convolution input layer and a linear output layer, and the so-called stacked cross-band and narrow-band blocks to build a mapping between complex-valued spectrums. The narrow-band block, composed of Mamba layers and convolutional layers along the time axis, processes each frequency independently with the same network parameters. The cross-band block, composed of convolutional layers along the frequency axis and full-band linear modules, processes each frame independently with the same network parameters. oSpatialNet-Mamba was initially proposed for online multi-channel applications. However, for single-channel applications, the spatial information still lies in RIR. The narrow-band block, designed based on CTF approximation, is able to learn such spatial information and thereby enabling single-channel speech dereverberation. Subsequent research has evaluated the effectiveness of oSpatialNet for single-channel speech enhancement [61].
In VINP, we adopt TCN+SA+S and a modified version oSpatialNet-Mamba, resulting in two different versions marked as ’VINP-TCN+SA+S’ and ’VINP-oSpatialNet’, respectively. For both versions, the input and output of DNNs are the 10-based logarithmic magnitude spectra. To avoid numerical issues, a small constant is added to the magnitude spectra before taking the logarithm. In VINP-TCN+SA+S, we use the same structure and settings as the original TCN+SA+S as shown in Fig. 3a, except that there is no activation function after the output layer, and we do not use dropout. In VINP-oSpatialNet, the DNN utilizes a 1-D convolution layer with a kernel size of to expand the input to 96 dimensions. The ’cross-band block’ and ’narrow-band block’ follow the original definition in oSpatialNet-Mamba, except that we replace the second forward Mamba layer in the narrow-band block with a backward Mamba layer by simply reversing the input and output along the time dimension for offline processing. The remaining modules are the same as those in the original oSpatialNet-Mamba, as depicted in Fig. 3b. We use VINP-TCN+SA+S and VINP-oSpatialNet to demonstrate the performance of two DNNs with different architectures and capabilities.
@@ IV-B3 Training Configuration
For DNN training, we set the hyperparameter in the loss function to . The speech utterances are segmented into 3 s. Each epoch contains 97,092 samples. The batch sizes of VINP+TCN+SA+S and VINP-oSpatialNet are set to 16 and 4, respectively. The AdamW optimizer [62] with an initial learning rate of 0.001 is employed. The learning rate exponentially decays with and in VINP-TCN+SA+S and VINP-oSpatialNet, respectively. Gradient clipping is applied with a L2-norm threshold of 10. The training is carried out for 800,000 steps in total. We average the model weights of the best three epochs as the final model.
@@ IV-B4 VBI Settings
The length of the CTF filter is set to . The fixed smoothing factor in Eq. (22) is set to 0.7 to obtain a stable result. Since the fundamental frequency of human speech is always higher than 85 Hz [63], we ignore the three lowest frequency bands and set their coefficients to zero to avoid the effect of extremely low SNR in these bands. Therefore, VBI processes a total of 254 frequency bands.
@@ IV-B5 Pseudo Excitation Signal
We use a logarithmic sine sweep signal with a frequency range of 62.5 Hz to 8000 Hz and a duration of 8.192 s as the pseudo excitation signal. Fade-in and fade-out with 256 and 128 samples are applied to the excitation signal to mitigate spectral leakage. The formulas for the excitation signal and the corresponding inverse filter can be found in Eq. (32) and [49, 50].
@@ IV-B6 RT60 and DRR Estimation
RT60 and DRR are key acoustic parameters characterizing the properties of RIR, and thus serve to evaluate the accuracy of RIR identification.
@@
Since sound energy decays exponentially over time, a linear fitting is applied to a segment of the logarithmic EDC, and the slope is used to calculate RT60. In this process, the key to linear fitting lies in the heuristic strategy for selecting the fitting range. In this work, the starting sampling point of the fitting is chosen within the range from 5 dB below the direct-path peak to the point corresponding to a 50 ms delay after the direct path. Meanwhile, the ending sampling point is defined as a point with 5 dB attenuation relative to the starting point. We fit all intervals that meet the conditions and adopt the result with the maximum absolute Pearson correlation coefficient for the RT60 estimation process as
@@
DRR refers to the ratio of the direct-path sound energy to the reverberant sound energy. In this work, the DRR is defined as
where the direct-path signal arrives at the th sample, and is the additional sample spread for the direct-path response, which typically corresponds to a 2.5 ms duration [59].
@@ IV-C1 Speech Dereverberation
We compare VINP with various advanced dereverberation approaches, including GWPE [7], SkipConvNet [65], CMGAN [26], and StoRM [28]. Additionally, comparisons are also made with the backbones TCN+SA+S [18] and the modified oSpatialNet-Mamba [60], which are trained using their original loss functions and input/output formats. In TCN+SA+S, we use the recommended Griffin-Lim’s iterative algorithm [66] to restore the phase spectrum. In the modified oSpatialNet-Mamba, we use the same DNN as in VINP-oSpatialNet and expand the number of channels in the input and output layers to 2 to process complex-valued spectra. We mark the modified DNN architecture as ’oSpatialNet*’. To further illustrate the impact of VEM iterations, we also show results of VINP-TCN+SA+S and VINP-oSpatialNet without VEM iterations (i.e., using only the DNNs in VINP to enhance the magnitude spectrum and keeping the phase spectrum reverberant, resulting in a spectrum as ). All comparison methods are trained and tested with their official codes (if available). All DL-based approaches are trained from scratch on the same training set. GWPE is implemented using the NaraWPE python package [67].
@@
VINP-TCN+SA+S (prop.) | 4.7 | 0.7+0.27iterations |
w/o VEM | 4.7 | 0.7 |
VINP-oSpatialNet (prop.) | 1.7 | 36.6+0.27iterations |
w/o VEM | 1.7 | 36.6 |
@@ IV-C2 Blind RIR Identification
We evaluate the effectiveness of RIR identification by measuring RT60 and DRR from the RIR estimates. The comparison methods include Jeub’s blind DRR estimation approach [68], FiNS [36] and BUDDy [38]. All these methods are implemented using the official codes (if available). Specifically, we modified the convolutional kernel size and STFT settings in FiNS to enable the network to process audio with a 16kHz sampling rate. As an unsupervised approach, we directly use the pretrained BUDDy model without retraining. For both FiNS and BUDDy, we employ the same implementation as in VINP to estimate RT60s and DRRs, as specified in Eq. (38) and Eq. (39).
Method | SimData | RealData | ||||||||||||
15 | ||||||||||||||
MOS | PESQ | ESTOI | LSD | DNSMOS | WER (%) | DNSMOS | WER (%) | |||||||
15 | ||||||||||||||
P.835 | P.808 | tiny | small | medium | P.835 | P.808 | tiny | small | medium | |||||
Unprocessed | 2.13 | 1.48 | 0.70 | 5.55 | 2.26 | 3.20 | 13.3 | 5.7 | 4.6 | 1.38 | 2.82 | 20.4 | 7.5 | 5.6 |
Clean | 4.32 | - | - | - | 3.28 | 3.90 | 7.4 | 4.5 | 4.0 | - | - | - | - | - |
GWPE [7] | - | 1.55 | 0.72 | 5.42 | 2.30 | 3.22 | 12.4 | 5.6 | 4.6 | 1.47 | 2.83 | 18.4 | 6.9 | 5.5 |
SkipConvNet [65] | - | 2.12 | 0.81 | 3.19 | 2.91 | 3.60 | 13.3 | 6.3 | 5.2 | 2.66 | 3.40 | 22.2 | 9.5 | 7.3 |
CMGAN [26] | 3.27 | 2.80 | 0.90 | 3.86 | 3.34 | 3.79 | 9.6 | 5.2 | 4.4 | 3.36 | 3.99 | 10.6 | 5.5 | 4.6 |
StoRM [28] | 3.97 | 2.36 | 0.86 | 3.59 | 3.24 | 3.93 | 11.0 | 6.1 | 5.1 | 3.20 | 3.94 | 17.0 | 9.2 | 7.2 |
TCN+SA+S [18] | 2.97 | 2.60 | 0.86 | 3.70 | 3.12 | 3.73 | 12.2 | 6.6 | 5.5 | 3.03 | 3.73 | 21.6 | 12.8 | 10.2 |
oSpatialNet* [60] | 3.71 | 2.87 | 0.92 | 3.03 | 3.16 | 3.88 | 9.1 | 5.0 | 4.2 | 3.10 | 3.87 | 10.5 | 5.5 | 4.6 |
VINP-TCN+SA+S (prop.) | 4.01 | 2.52 | 0.87 | 3.25 | 3.10 | 3.88 | 9.0 | 5.1 | 4.4 | 2.89 | 3.77 | 11.8 | 6.1 | 5.2 |
w/o VEM | - | 2.27 | 0.84 | 3.59 | 3.09 | 3.67 | 11.7 | 6.2 | 5.2 | 2.86 | 3.37 | 17.0 | 8.4 | 7.1 |
VINP-oSpatialNet (prop.) | 4.17 | 2.82 | 0.90 | 3.17 | 3.11 | 3.86 | 8.2 | 4.9 | 4.2 | 3.00 | 3.80 | 9.1 | 5.1 | 4.4 |
w/o VEM | - | 2.69 | 0.90 | 3.33 | 2.88 | 3.74 | 8.3 | 5.0 | 4.2 | 2.98 | 3.58 | 9.6 | 5.2 | 4.5 |